4. The total cost (TC) and demand equations for a monopolist is

TC=100+5Q2

P=200-5Q



a)     What is the profit-maximizing quantity?

b)     What is the profit-maximizing price?


1
Expert's answer
2021-06-23T12:47:56-0400

"TC=100+5Q^2"

 "MC = \frac{dTC}{dQ}"

"= 0+10Q"

"Monopolistic maximizing condition""MR = MC"

"P = 200-5Q"

"TR = PQ"

"(200-5Q)Q"

"= 200Q-5Q^2"

"MR = \frac{dTR}{dQ}"

"MR = 200-10Q"

"MR = MC"

"200-10Q = 10Q"

"200 = 10Q +10Q"

"200= 20Q"

"Q = \frac{200}{20}"

"Q = 10"

"Substituting into P equation"

"P = 200-5Q"

"P = 200-5(10)"

"P = 200-50"

"Profit-Maximizing price = 150""b) Profit-Maximizing quantity""P = MC"

"200-5Q = 10Q"

"200 = 10Q + 5Q"

"200 = 15Q"

"Q = \frac{200}{15}"

"Profit-Maximizing Quantity = 13.3"